OpenStudy (anonymous):

I really need help with this question ;_; What is the missing exponent? (10^6) = 10 ^-18

Nnesha (nnesha):

\[\huge\rm (x^m)^n = x^{m \times n}\] let's say missing exponent is x \[\huge\rm (10^6)^x =10^{-18}\]

Nnesha (nnesha):

multiply the exponents let's say missing exponent is x \[\huge\rm \huge\rm (10^6)^x =10^{-18}\] \[\huge\rm 10^{6 \times x}=10^{-18}\] bases are the same so you cancel them ot then solve for x in other words what would you multiply with 6 yo get -18

Nnesha (nnesha):

to get*

OpenStudy (anonymous):

-3 @Nnesha

Nnesha (nnesha):

yep

OpenStudy (anonymous):

Thank you! <3

Nnesha (nnesha):

\[\huge\rm \cancel{10}^{6 \times x}=\cancel{10}^{-18}\] \[6x=-18\] solve for x divide both sides by 6 you will get the answer

Nnesha (nnesha):

my pleasure :=) welcome to open study!

OpenStudy (anonymous):

Thanks ^-^

Nnesha (nnesha):

o^_^oyw